OpenStudy (anonymous):

Thalia has 65 red balloons and 52 blue balloons. What is the greatest common factor of 65 and 52? F. 15 G. 13 H. 5 J. 4 sister's question

OpenStudy (sayan):

G. 65=5*13 52=2*2*13 so there is 13 exist in both. Good luck

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Thanks for the medal!!!!